OpenStudy (anonymous):

f(x)= 13/x-11+ 1/x-9 ? List all real values of x such that f(x)=0

OpenStudy (anonymous):

\[f(x)=\frac{13}{x-11}+\frac{1}{x-9}\]?

OpenStudy (anonymous):

if so set \[\frac{13}{x-11}+\frac{1}{x-9}=0\] \[13(x-9)+(x-11)=0\] \[14 x-128 = 0\] \[x=\frac{128}{14}=\frac{64}{7}\]

OpenStudy (anonymous):

Thanks satellite73!!